# Editing Factual Knowledge and Explanatory Ability of Medical Large Language Models

### v2

@@ 4. Experiments
Figure 4. Knowledge type distribution of MedFE dataset.
@@ 4.1. Experimental Setup
we utilize the metrics constructed in Section 3.4 in our evaluation. The computation of other metrics follows EasyEdit (Wang et al., 2023a), which are measured as average accuracy between the token matching of the predicted output and the expected output. For ease of presentation, we employ abbreviations to represent each metric: Efficacy (Eff.), Generality (Gen.), Locality (Loc.), and Fluency (Flu.). Due to limitations of original data, we measure TD, EM, SS, TS for MedCF and measure TS and CT for MedFE. To examine the trade-off between edit success and locality, we further report the weighted mean by:
@@
As there is a lack of datasets for medical model editing, all our experiments are conducted on MedCF and MedFE datasets, using four NVIDIA V100 32G. In our experiments, we specifically address the single-edit problem, which has also been adopted in previous studies (Meng et al., 2022a; Dai et al., 2022; Huang et al., 2023). To tackle this problem, we utilize two medical LLMs, namely ChatDoctor-13B (Yunxiang et al., 2023) and Meditron-7B (Chen et al., 2023), which serve as the to-be-edited models in our research. We compare various model editing methods, including FT (Fine-tuning on multiple layers), LoRA (Hu et al., 2021), ROME (Meng et al., 2022a), MEND (Mitchell et al., 2022a), and MEMIT (Meng et al., 2022b). All hyper-parameters of baselines are set according to optimal values in the validation set of corresponding works. During the hyper-parameters selection of MedLaSA, we search for $r$, $\alpha$, and editable weights $W$. The search scope for $r$ and $\alpha$ was limited to the values [2, 8, 24, 32, 64, 128]. The analysis of $r$ and $\alpha$ can be seen in Figure 5. The weights that were editable in ChatDoctor and Meditron include Attn weights $W_{q}$, $W_{v}$, $W_{k}$, $W_{o}$, and MLP weights $W_{up}$, $W_{down}$, $W_{gate}$, which is analysed in Table 7. We selected parameter with the highest average value in the validation set. The hyper-parameters we use are as follows:
  • ChatDoctor-13B on MedCF dataset: $r$: 24, $\alpha$: 32, learning rate: 1e-05, the number of steps: 70, the maximum length: 40, the editable weights: $W_{q},W_{v},W_{up},W_{down}$.
  • Meditron-7B on MedCF dataset: $r$: 24, $\alpha$: 32, learning rate: 1e-05, the number of steps: 70, the maximum length: 40, the editable weights: $W_{q},W_{v},W_{k},W_{o},W_{up},W_{down},W_{gate}$.
  • ChatDoctor-13B on MedFE dataset: $r$: 24, $\alpha$: 32, learning rate: 5e-05, the number of steps: 40, the maximum length: 200, the editable weights: $W_{up},W_{down}$.
  • Meditron-7B on MedFE dataset: $r$: 24, $\alpha$: 24, learning rate: 5e-05, the number of steps: 40, the maximum length: 200, the editable weights: $W_{q},W_{v},W_{k},W_{o},W_{up},W_{down},W_{gate}$.
@@
Table 3. Statistics of MedCF and MedFE. The number of samples for different types is kept consistent, further details are provided in the Figure 3 and 4.
@@ 4.2. Main Results (RQ1)
We present the main results compared with baselines. As indicated in Table 4, MedLaSA demonstrates significant improvements over baselines across most metrics. For instance, MedLaSA exhibits superior performance of Fluency on both datasets and two LLM backbones, validating our method’s ability to maintain generation capability. The experimental results of FT indicate that excessive retraining of parameters could result in model collapse, causing the model to lose its original generating capability (i.e., much lower Fluency). LoRA introduces supplementary parameters but fails to consider the impact on unrelated knowledge. It also overlooks the specific positioning of knowledge in LLMs and dependency of knowledge on different layers. As a result, when compared to MedLaSA, LoRA may have a similar level of editing success, but it performs poorly in terms of Locality. MEND has high requirements for initialization conditions, resulting in lower average performance on these medical datasets. ROME focuses solely on single-layer knowledge editing of LLMs, without taking into account the knowledge stored in multi-layers, thus the performance tends to deteriorate.
MEMIT performs well in terms of Efficacy and Generality on MedCF because it is specifically designed to handle counterfactual data based on triplets, similar to MedCF. It edits counterfactual data by locating the key through the subject in the prompt and optimizing the value to select the object. However, MEMIT falls short in terms of Locality when compared to MedLaSA, particularly in Entity Mapping (EM value decreased by 33.12% on ChatDoctor). This is because MEMIT only learns the mapping relationship between the head and tail entities, leading to consistent predictions when the subject in the locality prompt is the same as editing prompt. Furthermore, MEMIT’s performance on the MedFE dataset is inferior due to its inability to handle long text output and complex multiple knowledge. Due to its reliance on subject-to-object localization, MEMIT may not be suitable for scenarios that involve complex knowledge and longer text. In contrast, our proposed MedLaSA addresses this issue by dynamically adjusting the scale of additional parameters and ensuring the insertion of complex knowledge.
@@ 4.4. Ablation Study (RQ3)
In this section, we analyze the effects on the performance of the model after removing Scaling $\alpha$ (SA) and Scaling Rank (SR) in the self-attention (Attn) and MLP layers, as shown in Table 6. We observe that removing SR leads to a decline in all metrics, with Locality-TS experiencing a decrease of approximately 5%. This suggests that SR plays a crucial role in maintaining the overall performance, especially on the Locality metrics. On the other hand, when SA is removed, there is an improvement in the Generality. This improvement, however, comes at the cost of significant decreases in Locality-CT (from 89.25% to 82.16%) and Locality-TS (from 84.11% to 75.84%), indicating that while SA helps in minimizing the model’s modification of irrelevant knowledge, it concurrently compromises the model’s generalization to rephrase. When both SA and SR are removed, the overall performance declines the most. It is worth noting that there are significant decreases in the Efficacy and Generality metrics, showing a decline in the success rate of medical model editing. Similar results can also be observed when our proposed method is applied exclusively to Attn weights or MLP weights, which further demonstrates the effectiveness of the proposed parameters SA and SR in medical model editing.
@@
Table 6. Ablation study on MedFE. The abbreviation ‘w/o’ indicates that the following module is removed, and the term ‘ALL’ indicates that both scaling strategies are employed in both Attn and MLP layers.
@@ 4.5. Hyper-parameters Analysis (RQ4)
Figure 5. Analysis of hyper-parameters $r_{o}$ and $\alpha_{o}$. The top two figures depict the results on MedFE, while the bottom two figures display the results on MedCF.
@@ 4.6. Comparison of Editable Weights (RQ5)
To qualitatively measure the significance of different editable weights and analyze the specific combination of weights that are more suitable for editing in MedLaSA, we conduct a comparison of editing weights in ChatDoctor on the MedFE dataset, as shown in Table 7. It can be seen that editing Attn weights $W_{q}$ and $W_{v}$ demonstrate a higher editing success rate compared to Attn weights $W_{k}$ and $W_{o}$, and it exhibits better generalization ability for rephrasing text. However, editing $W_{q}$ and $W_{v}$ shows a poorer performance in terms of the Locality metrics. When all four weights are edited together, the learning of editing text is improved, but this improvement is accompanied by a decrease in the Locality metrics (e.g., from 98.48% to 90.23% on CT and from 97.82% to 86.21% on TS).
Figure 6. Case of causal tracing on MedCF dataset. The presented red/green heatmaps illustrate the impact of Attn/MLP weight restoration after corrupted input.
When comparing weights of MLP, $W_{up}$ consistently outperforms $W_{down}$ in all metrics. This suggests that $W_{up}$ may have ability to retain more knowledge and is more suitable for editing medical models. Furthermore, when $W_{up}$, $W_{down}$, and $W_{gate}$ are used together, there is a notable enhancement in both Efficacy and Average performance. Moreover, editing MLP weights ($W_{up}$, $W_{down}$, $W_{gate}$) typically leads to significantly higher editing success rate (Efficacy ¿ 80%). This aligns with the viewpoint expressed in previous papers that MLP is the primary storage for modeling world knowledge, and editing MLP layer yields better results compared to the attention layer (Meng et al., 2022a). It is worth noting that when both Attn and MLP weights are simultaneously made trainable for editing, there are additional enhancements in Efficacy and Generality. However, this comes with significant decreases in Locality and Fluency. This suggests that incorporating more trainable adapter parameters increases the success rate of medical model editing. Consequently, it leads to a stronger impact on the prediction of irrelevant information.

### v3

@@ 4. Experiments
Figure 3. Knowledge type distribution of MedCF (Left) and MedFE (Right) dataset.
@@ 4.1. Experimental Setup
we utilize the metrics constructed in Section 3.4 in our evaluation. The computation of metrics follows EasyEdit (Wang et al., 2023a), which are measured as average accuracy between the token matching of the predicted output and the expected output. For ease of presentation, we employ abbreviations to represent each metric: Efficacy (Eff.), Generality (Gen.), Locality (Loc.), and Fluency (Flu.). Due to limitations of original data, we measure TD, EM, SS, TS for MedCF and measure TS and CT for MedFE. To examine the trade-off between edit success and locality, we further report the weighted mean by:
@@
As there is a lack of datasets for medical model editing, all our experiments are conducted on MedCF and MedFE datasets, using four NVIDIA V100 32G. In our experiments, we specifically address the single-edit problem, which has also been adopted in previous studies (Meng et al., 2022a; Dai et al., 2022; Huang et al., 2023). To tackle this problem, we utilize two medical LLMs, namely ChatDoctor-7B (Yunxiang et al., 2023) and Meditron-7B (Chen et al., 2023), which serve as the to-be-edited models in our research. We compare various model editing methods, including FT (Fine-tuning on multiple layers), LoRA (Hu et al., 2021), ROME (Meng et al., 2022a), MEND (Mitchell et al., 2022a), and MEMIT (Meng et al., 2022b). The implementation of all baseline codes is based on the open-source tool EasyEdit (Wang et al., 2023a). We maintain same settings across all methods to ensure consistency. All hyper-parameters of baselines are set according to optimal values in the validation set of corresponding works. During the hyper-parameters selection of MedLaSA, we search for $r$, $\alpha$, and editable weights $W$. The search scope for $r$ and $\alpha$ was limited to the values [2, 8, 24, 32, 64, 128]. The analysis of $r$ and $\alpha$ can be seen in Figure 4. The weights that were editable in ChatDoctor and Meditron include Attn weights $W_{q}$, $W_{v}$, $W_{k}$, $W_{o}$, and MLP weights $W_{up}$, $W_{down}$, $W_{gate}$, which is analysed in Table 7. We selected parameter with the highest average value in the validation set.
@@
Table 3. Statistics of MedCF and MedFE.
@@ 4.2. Main Results (RQ1)
We present the main results compared with baselines. As indicated in Table 4, MedLaSA demonstrates significant improvements over baselines across most metrics. For instance, MedLaSA exhibits superior performance of Fluency on both datasets and two LLM backbones, validating our method’s ability to maintain generation capability. The experimental results of FT indicate that excessive retraining of parameters could result in model collapse, causing the model to lose its original generating capability (i.e., much lower Fluency). LoRA introduces supplementary parameters but fails to consider the impact on unrelated knowledge. It also overlooks the specific positioning of knowledge in LLMs and dependency of knowledge on different layers. As a result, LoRA performs poorly in terms of Locality compared to MedLaSA. MEND has high requirements for initialization conditions, resulting in lower average performance on these medical datasets. ROME focuses solely on single-layer knowledge editing of LLMs, without taking into account the knowledge stored in multi-layers, thus the performance tends to deteriorate.
MEMIT performs well in terms of Efficacy and Generality on MedCF because it is specifically designed to handle counterfactual data based on triplets, similar to MedCF. It edits counterfactual data by locating the key through the subject in the prompt and optimizing the value to select the object. This phenomenon is also evident in the case of ROME, as it has enabled ROME to achieve comparable or even superior performance to MedLaSA. However, MEMIT and ROME fall short in terms of Locality when compared to MedLaSA, particularly in Entity Mapping (EM value of MEMIT decreased by 33.12% on ChatDoctor). This is because MEMIT and ROME only learn the mapping relationship between the head and tail entities, leading to consistent predictions when the subject in the locality prompt is the same as the editing prompt. Furthermore, MEMIT’s performance on the MedFE dataset is inferior due to its inability to handle long text output and complex multiple knowledge. Due to its reliance on subject-to-object localization, MEMIT may not be suitable for scenarios that involve complex knowledge and longer text. In contrast, our proposed MedLaSA addresses this issue by dynamically adjusting the scale of additional parameters and ensuring the insertion of complex knowledge.
@@ 4.4. Ablation Study (RQ3)
In this section, we analyze the effects after removing Scaling $\alpha$ (SA) and Scaling Rank (SR) in the self-attention (Attn) and MLP layers, as shown in Table 6. We observe that removing SR leads to a decline in all metrics. This suggests that SR plays a crucial role in maintaining the overall performance, especially on the Locality metrics. On the other hand, when SA is removed, there is an improvement in the Generality. This improvement, however, comes at the cost of significant decreases in Locality-CT (from 89.25% to 82.16%) and Locality-TS (from 84.11% to 75.84%), indicating that while SA helps in minimizing the model’s modification of irrelevant knowledge, it concurrently compromises the model’s generalization to rephrase. When both SA and SR are removed, the overall performance declines the most. It is worth noting that there are significant decreases in the Efficacy and Generality metrics, showing a decline in the success rate of medical model editing. Similar results can also be observed when our proposed method is applied exclusively to Attn weights or MLP weights, which further demonstrates the effectiveness of the proposed parameters SA and SR in medical model editing.
@@
Table 6. Ablation study on MedFE.
@@ 4.5. Hyper-parameters Analysis (RQ4)
Figure 4. Analysis of hyper-parameters $r_{o}$ and $\alpha_{o}$.
@@ 4.6. Comparison of Editable Weights (RQ5)
To qualitatively measure the significance of different editable weights and analyze the specific combination of weights that are more suitable for editing in MedLaSA, we conduct a comparison of editing weights in ChatDoctor on the MedFE dataset, as shown in Table 7. For Attn weights, when all four weights are edited together, the learning of editing text is improved, but Locality metrics decrease (e.g., from 98.48% to 90.23% on CT and from 97.82% to 86.21% on TS). When comparing weights of MLP, $W_{up}$ consistently outperforms $W_{down}$ in all metrics. This suggests that $W_{up}$ may have ability to retain more knowledge and is more suitable for editing medical models. Moreover, editing MLP weights ($W_{up}$, $W_{down}$, $W_{gate}$) typically leads to significantly higher editing success rate (Efficacy ¿ 80%). It is worth noting that when both Attn and MLP weights are simultaneously made trainable for editing, there are additional enhancements in Efficacy and Generality. However, this comes with significant decreases in Locality and Fluency. This suggests that incorporating more trainable adapter parameters increases the success rate of medical model editing. Consequently, it leads to a stronger impact on irrelevant information.
Figure 5. Cases of causal tracing on MedCF. The presented red/green heatmaps illustrate the impact of Attn/MLP weight restoration after corrupting input.
@@ 5. Conclusion
6. ACKNOWLEDGMENTS
This research was partially supported by CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program, Research Impact Fund (No.R1015-23), CityU - HKIDS Early Career Research Grant (No.9360163), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), and SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046). Additionally, this work was supported in part by the grants from National Natural Science Foundation of China (No.62222213, U23A20319, 62072423).